OpenStudy (anonymous):

Simplify r^2a / r^2a-3. 2a and 2a-3 are radicals. Can you explain, also? Thanks!

OpenStudy (anonymous):

\[\frac {r^2\sqrt{2a}} {r^2\sqrt{a-3}}\] Yes?

OpenStudy (anonymous):

Where do the radicals come from? Confused.. I don't think I learned this.

OpenStudy (anonymous):

Is that the correct equation? Your post says r^2a but then says 2a is a radical which does not make sense.

OpenStudy (anonymous):

\[r^{\sqrt{2a}}\] Or this?

OpenStudy (anonymous):

Oh, I'm sorry. I meant 2a and 2a-3 are exponents. Apologies.

OpenStudy (anonymous):

\[\frac{r^{2a}}{ r^{2a-3}}\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

\[r ^{2a} / r ^{2a-3}\]

OpenStudy (anonymous):

Yeah, that's right. (:

OpenStudy (anonymous):

Apply rule for exponents a^p /a^q = a^(p-q).

OpenStudy (anonymous):

So it's r^(2a - 2a-3)?

OpenStudy (anonymous):

yes, except it's 2a - (2a-3) which is different...

OpenStudy (anonymous):

Oh, yeah. Sorry. Thank you very much. (:

OpenStudy (anonymous):

u r welcome, if u don't like the equation editor, just use brackets to make things clear.